OpenStudy (anonymous):

The following table shows the probability distribution for a discrete random variable. X 22 24 25 30 33 37 P(X) 0.16 0.09 0.18 0.12 0.24 0.21 What is the mean of this discrete random variable? That is, what is E(X), the expected value of X? A. 27.5 B. 28.5 C. 29.47 D. 30.27

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

boring exercise, you have to multiply and add

OpenStudy (misty1212):

\[22\times .16+24\times .09+...\] all the way down the line

OpenStudy (amistre64):

so we have an X vector, and P(X) vector, and we want the dotproduct ...

OpenStudy (misty1212):

lol dot product !!

OpenStudy (amistre64):

X 22 24 25 30 33 37 P(X) 0.16 0.09 0.18 0.12 0.24 0.21 out of 100, we expect 16 values of 22 9 values of 24 18 values of 25 12 values of 30 24 values of 33 21 values of 37 sum of all events = +(22+22+...+22) ; 16 times +(24+24+...+24) ; 9 times +(25+25+...+25) ; 18 times +(30+30+...+30) ; 12 times +(33+33+...+33) ; 24 times +(37+37+...+37) ; 21 times ------------- n = 100 events mean is defined as ... sum of all events, divided by the number of events.

OpenStudy (misty1212):

or you could work with the decimals and not divide by 100 at the end right?

OpenStudy (amistre64):

just making a point is all, showing how we go from the definition of a mean to the sum of xP(x) \[\frac{22(16)+24(9)+25(18)+30(12)+33(24)+37(21)}{100}\] \[22(\frac{16}{100})+24(\frac{9}{100})+25(\frac{18}{100})+30(\frac{12}{100})+33(\frac{24}{100})+37(\frac{21}{100})\] \[22(.16)+24(.09)+25(.18)+30(.12)+33(.24)+37(2.1)\]

OpenStudy (amistre64):

.21 on the end :)

OpenStudy (anonymous):

so whats the answer?

OpenStudy (amistre64):

you want us to add and multiply for you? nah, thats on you.

OpenStudy (anonymous):

I was just confused...sorry. So, i add 22(.16)+24(.09)+25(.18)+30(.12)+33(.24)+37(2.1) all that and i multiply what?

OpenStudy (anonymous):

when i added everything i got 99.4

OpenStudy (amistre64):

lets see what the wolf gives us :) so we avoid any mistakes

OpenStudy (amistre64):

.21, not 2.1 22(.16)+24(.09)+25(.18)+30(.12)+33(.24)+37(.21) ^^^^ http://www.wolframalpha.com/input/?i=22%28.16%29%2B24%28.09%29%2B25%28.18%29%2B30%28.12%29%2B33%28.24%29%2B37%28.21%29 29.47

OpenStudy (anonymous):

ohhh okay thank you :-)

OpenStudy (amistre64):

yep